OpenStudy (anonymous):

A crystallographer measures the horizontal spacing between molecules in a crystal. The spacing is 11.75nm . What is the total width of a crystal in millimeters 10^5 molecules across?

OpenStudy (aaronq):

\(\sf 1~nm=1.0*10^{-9}m=1.0*10^{-6}mm\)